### v1

Abstract
Modern moderation systems increasingly support multiple languages, but often fail to address localisation and low-resource variants - creating safety gaps in real-world deployments. Small models offer a potential alternative to large LLMs, yet still demand considerable data and compute. We present LionGuard 2, a lightweight, multilingual moderation classifier tailored to the Singapore context, supporting English, Chinese, Malay, and partial Tamil. Built on pre-trained OpenAI embeddings and a multi-head ordinal classifier, LionGuard 2 outperforms several commercial and open-source systems across 17 benchmarks, including both Singapore-specific and public English datasets. The system is actively deployed within the Singapore Government, demonstrating practical efficacy at scale. Our findings show that high-quality local data and robust multilingual embeddings can achieve strong moderation performance, without fine-tuning large models. We release our model weights [^1] https://huggingface.co/govtech/lionguard-2 and part of our training data [^2] https://huggingface.co/datasets/govtech/lionguard-2-synthetic-instruct to support future work on LLM safety.
LionGuard 2: Building Lightweight, Data-Efficient & Localised Multilingual Content Moderators
Leanne Tan1thanks: Contributed equally, Gabriel Chua [^1] footnotemark: 1, Ziyu Ge2, Roy Ka-Wei Lee1,2, 1GovTech, Singapore, 2Singapore University of Technology and Design {leanne_tan|gabriel_chua}@tech.gov.sg Warning: this paper contains references and data that may be offensive.
1 Introduction
As AI systems are increasingly deployed across diverse linguistic communities, moderation systems [^3] We use the terms ”moderation system”, ”moderation classifier”, and ”guardrail” interchangeably to refer to text filters that assess content safety. are evolving to offer broader multilingual support OpenAI (2024); Meta (2025). However, their effectiveness in localised, low-resource, or code-mixed settings remains limited - leaving critical safety gaps. For instance, multilingual adversarial prompts have been shown to bypass robust filters Yong et al. (2024); Shen et al. (2024); Wang et al. (2024). Singapore exemplifies the challenge: everyday discourse blends English, Chinese, Malay, and Tamil in code-mixed forms like Singlish, featuring local slang, abbreviations, and dialectal variants (Figure 1). Moderation systems that ignore such linguistic nuance risk degraded performance and exploitation in real-world deployments.
Figure 1: Example of code-mixed Singlish combining English, Chinese, and Malay.
In this work, we present LionGuard 2, a lightweight, data-efficient moderation classifier tailored to Singapore’s multilingual context. Unlike large LLM-based guardrails Inan et al. (2023), LionGuard 2 uses pre-trained multilingual embeddings and a compact classifier head to achieve fast, scalable deployment with minimal compute and training data. This upgrade over our prior system, LionGuard 1 Foo and Khoo (2025), includes: (i) a richer risk taxonomy with severity levels, (ii) improved robustness to noisy and code-mixed input, and (iii) multilingual support across English, Chinese, Malay, and partial Tamil. LionGuard 2 can be retrained in under two minutes, runs on CPUs, and integrates easily into production workflows. We benchmark LionGuard 2 across 17 datasets, spanning both Singapore-localised and public English testbeds, and find that it outperforms both commercial and open-source moderation systems in F1 score.
Our contributions are threefold: (i) We present LionGuard 2 as a case study for building practical, localised moderation systems under resource constraints. (ii) We share empirical insights from model architecture choices, data curation strategies, and comparative evaluations. (iii) We release the classifier weights and a portion of our training data to support future research in LLM safety.
2 Related Work
2.1 Multilingual Content Moderation
While commercial moderation APIs offer multilingual support, their efficacy on low-resource or code-mixed languages is often unclear. Open-source models such as LlamaGuard Meta (2024b, 2025), DuoGuard Deng et al. (2025), and PolyGuard Kumar et al. (2025) provide broader coverage, but do not address cultural localisation, limiting their robustness in real-world multilingual environments.
Recent benchmarks Ng et al. (2024); Gupta et al. (2024); Chua et al. (2025) address this gap by evaluating models on Singapore-specific, code-mixed input. Our earlier system, LionGuard 1 Foo and Khoo (2025), showed that a multilingual embedding-based classifier can outperform LLM-based solutions on such tasks. However, it used a coarser risk taxonomy and lacked partial Tamil coverage. In this work, we present LionGuard 2, which improves performance on local and general benchmarks, introduces severity-aware ordinal heads, and extends multilingual robustness while maintaining a lightweight architecture.
2.2 Small, Inference-Efficient Guardrails
Recent work has trended toward smaller moderation classifiers. For example, LlamaGuard 3 (1B) Fedorov et al. (2024) and ShieldGemma (2B) Zeng et al. (2024) exemplify compact models designed for efficient inference. Other open-source efforts Kumar et al. (2025); Deng et al. (2025) also fine-tune small 0.5B and 2.5B models. However, training these models is costly, requiring large labeled datasets (often over a million examples) and substantial compute. This discourages efforts to customise guardrails for local safety contexts and ultimately limits adoption of safe AI deployments.
Conversely, embedding-based methods offer a complementary path. Systems can effectively achieve strong performance with pre-trained embeddings in retrieval and classification tasks Enevoldsen et al. (2025); Chen et al. (2024); Sturua et al. (2024), and are even available in specialised domains Tang and Yang (2025); VoyageAI (2024), proving its practicality.
3 System Overview
Figure 2: Example of LionGuard 2 working as a bidirectional filter around an LLM Chatbot.
LionGuard 2 is designed as a lightweight moderation system for any text content. Figure 2 illustrates an example of its role as both an input filter (screening user prompts) and an output filter (verifying model responses) before the text reaches the language model or end users respectively. LionGuard 2 can also be used in AI application safety testing, to detect if application responses contain unsafe elements.
Figure 6 demonstrates LionGuard 2 acting as an chatbot guardrail. LionGuard 2 flags localised unsafe content that bypasses both the LLM’s internal safety alignment and commercial moderation classifiers (i.e. OpenAI Moderation). In this example, Singapore-specific acronyms and slang are used to elicit unsafe content from the LLM.
Real World Deployment.
LionGuard 2 replaces its predecessor and is deployed on the Singapore Government’s AI Guardian platform GovTech Singapore (2025b) as a safety module for any text-centric service requiring localised safeguards. Developers can easily apply LionGuard within the standard Chat Completions API request GovTech Singapore (2025a).
Running synchronously on a single CPU, the embedding call handles $\approx 250\;\text{tokens}\,\text{s}^{-1}$, while the classifier head itself processes $\approx 1.5\times 10^{4}\;\text{tokens}\,\text{s}^{-1}$, giving an end-to-end throughput of $\approx 300\;\text{tokens}\,\text{s}^{-1}$. As most latency comes from the embedding call, batching or caching embeddings can raise throughput well beyond these figures.
4 Methodology
Figure 3: The LionGuard 2 methodology. We consolidate and process multiple data sources, apply semi-supervised labeling with human-aligned LLMs, and train a lightweight classifier on embeddings from a carefully chosen model.
4.1 Data Curation
4.1.1 Safety Taxonomy
Every content-moderation system defines its own harm categories, and we adopt the two-level taxonomy in Appendix B, originally proposed by Goh et al. (2025) and adopted in Chua et al. (2025) for the Singapore context. While this taxonomy is customised to our organisational needs, each label can be mapped to other major frameworks by MLCommons Vidgen et al. (2024a), OpenAI OpenAI (2024), and the major cloud providers Azure ; AWS (2025). All subsequent LionGuard 2 design choices are aligned with this internal and localised taxonomy.
We encourage practitioners adopting similar methodologies to begin their projects with a comprehensive, effective taxonomy that matches their real-world use case.
4.1.2 Data Sources.
Our goal is to curate a small yet rich set of texts that reflects Singaporean discourse. We first combined three data sources:
  1. 1.
    Local comments. We extract texts from Singaporean forums and subreddits, previously described in Foo and Khoo (2025).
  2. 2.
    Synthetic queries. To broaden style coverage, each local comment was rewritten by gpt-4o-mini into a chatbot query, then verified and refined using self-reflection and chain-of-thought (CoT) prompting. (see prompt in Appendix C.1). Figure 4 illustrates an example of an transformed comment.
  3. 3.
    Open-source english data. Open-source english datasets containing text relevant to our safety taxonomy were added (See Appendix C.2). Some datasets were eventually excluded after ablation revealed binary $F_{1}$ loss of as much as 30%.
Figure 4: Example of a synthetically augmented Singaporean comment.
Initial experiments also included using gpt-4o-mini to translate local comments into Malay and Tamil; however, these variants lowered downstream $F_{1}$ and were removed (Appendix C.3). From our manual review, we hypothesise that the LLM struggles to translate toxic content across languages, either losing toxicity or failing to fully grasp the context. Notably, our final corpus contains no machine-translated Chinese, Malay or Tamil, and all non-English text appear only in naturally code-mixed Singlish.
4.1.3 Automated Labelling with Human Supervision
To mimic low-resource constraints, we employ LLMs to annotate the training examples, but employ statistical methods to ensure as much alignment as possible with human labellers. We begin with a panel of six humans and six LLMs 444o3-mini-low OpenAI (2025), Gemini 2.0 Flash Google (2025), Claude 3.5 Haiku Anthropic (2024), Llama 3.3 70B Meta (2024a), Mistral Small 3 Mistral (2025), and AWS Nova Lite Intelligence (2024). Through the Alt-Test methodology (Calderon et al., 2025), we identified Gemini 2.0 Flash, o3-mini-low, and Claude 3.5 Haiku to be best aligned with our six human annotators, and labeled every example in our dataset with these three selected models (system prompt in Appendix C.5).
4.1.4 Data Filtering.
A five-stage funnel data (Figure 3) pipeline was used in data curation, systematically adjusting parameters (Appendix C.6) at each stage.
Resulting Corpus.
The final training set contained 26,207 unique texts: 20,333 online comments, 2,098 synthetically augmented comments, and 3,776 texts from open-source English datasets. The test set contained 6,249 raw comments and 7,058 synthetic comments. [^5] In this study, we did not adopt the original train-test split from Foo and Khoo (2025), and instead used a time-series based split. Our training set is significantly smaller than that of other methods that fine-tune decoder-based models for content moderation, and is also 70% smaller than what was used for LionGuard 1.
4.2 Architecture
4.2.1 Selecting Domain-Specific Embeddings
Embedding choice is critical as it defines the representation space on which all downstream moderation classifier layers operate. We selected six multilingual open- and closed-source text embedding models and trained the same multi-head classifier on each set of embeddings.
Embeddings Test RabakBench
SS ZH MS TA
text-embedding-3-large $3,072_{d}$ OpenAI (2025) 77.0 88.1 87.8 78.4 66.6
cohere-embed-multilingual-v3.0 $1,024_{d}$ Cohere (2025a) 72.9 64.2 67.9 60.9 56.4
cohere-embed-v4.0 $1,536_{d}$ Cohere (2025b) 69.0 61.1 66.2 38.9 3.8
BGE-M3 $1,024_{d}$ Chen et al. (2024) 63.2 51.9 65.1 60.6 51.0
snowflake-arctic-embed-l-v2.0 $1,024_{d}$ Yu et al. (2024) 64.3 44.6 55.0 45.4 41.7
Qwen3-Embedding-0.6B $1,024_{d}$ Zhang et al. (2025) 87.2 61.4 67.9 60.9 56.4
Table 1: Binary $F_{1}$ when swapping sentence encoders.
Selection Outcome.
Results on our different hold out sets (Table 1) show that text-embedding-3-large achieved the highest binary $F_{1}$, scoring as much as 20% above the next-best model. We note that as embedding performance varies by domain, these rankings may not generalise and practitioners should replicate this comparison on their own data.
4.2.2 Training a Lightweight Classifier
The pre-trained embeddings are frozen and fed into a trainable multi-head network (Figure 7).
Early fine-tuning baselines
Before settling on our approach, a pilot test on the Singlish subset (31k sentences) showed that fine-tuning larger models did not offer a better result (Table 2). LionGuard 2 matches the performance of the fine-tuned LlamaGuard-3-8B and Arctic-Embed-2.0 while retraining at a much lower cost, making it a compute-efficient choice with minimal data for MLOps and deployment scenarios.
Model Test $RB_{SS}$ hw/time
OpenAI embeddings + 6 heads 82.9 85.8 CPU/60s
LlamaGuard-3-8B + LoRA 82.1 89.8 A100 40GB/16h
arctic-embed-l-v2.0 + 6 heads 74.1 67.3 T4 16GB/3h
Table 2: Binary $F_{1}$ for fine-tuned model variants, on our test set and RabakBench (Singlish); full configs in Appendix D.2.
Ordinal heads for Level-2 harms.
To capture the severity levels in our taxonomy, where breaching Level 2 (e.g., hate speech) would imply breaching Level 1 (e.g., discriminatory statements), the classification heads have a two-output design:
$(p_{1},\,p_{2})$ $=\sigma\!\bigl{(}\mathrm{Dense}_{2}(\mathbf{h})\bigr{)},$ (1)
$p_{1}$ $=P(y_{c}>0)$ (Level 1)
$p_{2}$ $=P(y_{c}>1)$ (Level 2)
[…]
In addition to the six category heads (one per risk category), we attached a single binary head (safe/unsafe) as we found it to consistently boost overall F1. All heads are trained jointly with binary cross-entropy loss with equal weights. We detail the training setup in Appendix D.3. The resulting classifier contains 0.85M parameters and occupies only 3.2 MB on disk.
Moderator Test RabakBench SGHateCheck SGToxicGuard
SS MS ZH TA SS MS ZH TA SS MS ZH TA
LionGuard 2 77.0 88.1 87.8 78.4 66.6 98.8 92.1 97.4 64.5 99.7 98.2 99.2 71.5
LionGuard 1.1 53.7 58.4 57.1 70.7 69.1 45.5 37.4 22.9 17.4 24.2 10.7 9.6 7.4
OpenAI Moderation 54.7 64.0 69.7 66.1 7.4 89.6 70.4 80.3 3.8 77.3 43.9 57.8 1.3
AWS Bedrock Guardrails 57.1 69.6 - 21.1 - 82.2 - 40.6 - 91.5 - 74.2 -
Azure AI Content Safety 53.8 66.0 67.0 66.2 48.7 76.8 67.9 68.4 75.3 69.5 54.3 63.2 30.1
Google Cloud Model Armor 36.8 62.5 68.3 74.0 73.4 81.8 74.0 89.3 63.5 83.3 82.8 88.8 71.9
LlamaGuard 3 8B 27.1 55.2 53.6 53.1 47.3 85.9 79.1 80.1 72.1 94.7 90.8 92.0 87.6
LlamaGuard 4 12B 26.5 60.6 54.6 65.2 73.0 68.8 57.4 63.9 58.8 78.6 74.3 77.0 77.9
Table 3: Binary $F_{1}$ scores on Singapore-localised benchmarks. The best results for each dataset are made bold. (–) indicates that the model does not support that language.
Moderator BT SRY-B OAI SST
LionGuard 2 73.7 73.7 70.5 100.0
LionGuard 1 35.0 31.6 55.8 34.7
OpenAI Mod 65.4 45.3 77.1 81.0
AWS Bedrock 76.4 50.7 77.4 84.4
Azure C. Safety 54.6 44.7 70.6 59.3
GCP Model Armor 51.3 42.7 74.8 46.3
LlamaGuard 3 8B 68.2 62.5 82.2 87.6
LlamaGuard 4 12B 67.0 58.7 77.5 98.0
Table 4: Binary $F_{1}$ scores on general English benchmarks - BeaverTails (BT), SORRY-Bench (SRY-B), OpenAI Moderation (OAI) and SimpleSafetyTests (SST). SST and SRY-B contain only unsafe prompts and thus the reported $F_{1}$ reflects recall.
5 Evaluation
5.1 Performance on 17 Benchmarks.
We compare LionGuard 2 against six moderation systems (OpenAI, 2024; AWS, 2025; Azure, ; GCP, 2025; Meta, 2024b, 2025) plus LionGuard 1 (Foo and Khoo, 2025) on 1 internal test set and 16 public benchmarks, including 13 localised datasets from Chua et al. (2025); Ng et al. (2024) and 4 general English datasets from Ji et al. (2023); Xie et al. (2025); Markov et al. (2023); Vidgen et al. (2024b).
Following prior moderation work Chi et al. (2024); Han et al. (2024), we report binary $F_{1}$ at a 0.5 threshold. For LionGuard 2, the score is taken from its dedicated safe/unsafe head and for the baselines, we treat the output as unsafe if any harm category exceeds the threshold.
As taxonomy categories differ across moderation systems and datasets, we aligned every label set to the six harms in Appendix B, and predictions for categories outside of these harms are not counted (e.g., Personal Identifiable Information, Medical Advice). Complete mappings are provided in Appendix E.
Results.
Table 3 reports binary $F_{1}$ across all benchmarks. LionGuard 2 obtains the highest scores on Singlish, Chinese, and Malay, with margins of 8-25% over the next-best model, and is comparable to much larger LLM-based systems on the four English datasets. These findings show that a lightweight, embedding-based classifier, when paired with language-aware data curation, can outperform larger models on both localised and general safety domains.
Category breakdown.
Full per-category results are listed in Appendix E.3. Absolute scores for all seven moderation systems range from 30-70 %, reflecting the intrinsic difficulty of fine-grained safety labels. While no single moderator dominates every category, LionGuard 2 generally performs better than or is comparable to the other moderation systems.
5.2 Robustness.
Practical moderation systems must handle noisy user input. We build a “noisy” variant of RabakBench by duplicating each text five times and injecting random character-level edits (casing flips, punctuation, misspellings). The results in Table 5 show that LionGuard 2 marginal 1.5% binary $F_{1}$ drop, effectively displaying tolerance to noise.
Moderator $RB_{SS}$ $RB_{SS\_noise}$
LionGuard 2 87.1 85.6
LionGuard 1 58.4 64.2
OpenAI Mod 64.0 52.2
Table 5: Binary $F_{1}$ on RabakBench Singlish with and without injected noise. LionGuard 2 remains robust, dropping only 1.5%.
5.3 Red-Teaming by Native Speakers
Employing native annotators, we curated a multilingual test set of 391 cases and benchmarked LionGuard 2 against five other safety models (see details in Appendix E.2). LionGuard 2 surpasses its nearest competitor by 2.4% in Chinese $F_{1}$ and 8.3% in Malay $F_{1}$. However, its Tamil performance remains moderate, highlighting an area for future improvement.
6 Key Insights
The observations gathered from our work are specific to LionGuard 2 and whether the same holds for other domains remains open. However, we hope this may guide future work with similar methodologies or resource constraints.
6.1 Localised data matters most
On the same architecture, training on authentic Singaporean comments alone achieved an average $F_{1}$ of 80.3%, whilst training on public English datasets alone achieved 45.3%. The combination bumped up results to an average $F_{1}$ of 81.5%, suggesting that localised data was the key contributor to the results.
6.2 Small models can outperform large models
In our fine-tuning experiments (Table 2), LlamaGuard-3-8B achieves similar test set performance and scores only 3% higher binary $F_{1}$ than LionGuard 2 on RabakBench. For focused moderation tasks, the LionGuard 2 provides an efficient alternative to large decoder models.
6.3 Embedding choice is decisive
OpenAI’s text embedding-3-large achieved the best F1 despite showing a similar or lower multilingual cosine alignment than cohere-embed-multilingual-v3.0 and BGE-M3 (Appendix D.1). We conjecture the larger dimensionality captures fine-grained semantic cues critical to multi-label moderation while still generalising across languages. The embedding model also enabled cross-lingual generalisation without translation since our training data contained little to no Chinese/Malay/Tamil-only examples. Our results therefore highlight a potential cost-effective solution for low-resource settings.
7 Limitations.
7.1 Reliance on closed-source embeddings
LionGuard 2 inherits its representations from OpenAI’s text-embedding-3-large. Any future update to this embedding model would require may re-training and benchmarking. Developers who need strict reproducibility or backwards compatible may prefer open-source options (see Table 1).
7.2 Misalignment between binary and category labels
About 4% of examples aggregated across two localised and three general datasets show disagreement between the binary head and category heads (Appendix E.1). Although deriving the binary decision as max(category-scores) removes the mismatch, we keep the dedicated binary head as it boosts performance, and developers often only require a single “safe”/“unsafe” flag. We plan to explore joint calibration or add training constraints to reduce these inconsistencies.
7.3 Lower performance for Tamil
All tested embedding models (including Tamil-centric sarvam-m) underperformed on Tamil, and adding LLM-translated Tamil data worsened results (Appendix C.3). Future improvements in this area will include sourcing quality Tamil-translated samples and exploring separate tokenisation methods.
8 Conclusion
LionGuard 2 is currently deployed across internal Singapore Government systems, validating that a lightweight classifier, built on strong multilingual embeddings and curated local data, can deliver robust performance in both localised and general moderation tasks. Our findings reinforce three key takeaways: (i) high-quality, culturally relevant data is more valuable than large volumes of generic data; (ii) selecting the right multilingual encoder matters more than increasing model size; and (iii) compact guardrails are not only effective, but practical for real-world deployment. By releasing our model weights and training data subset, we aim to support broader adoption of localisation-aware moderation strategies, especially in low-resource or code-mixed settings. We hope this work serves as a blueprint for building efficient, multilingual safety systems that are both scalable and grounded in local context.
Ethical Considerations
Potential Harms
While LionGuard 2 demonstrates effective performance in moderating localised unsafe content, we acknowledge that the system is not foolproof. Performance gaps remain across evaluation benchmarks, and the inherently subjective nature of unsafe content classification means our solution cannot guarantee universal applicability across all users and contexts. Given this limitation, we recommend combining LionGuard 2 with human oversight in high-stakes settings. Users should be aware of potential system failures and underperformance, particularly when dealing with edge cases or evolving harmful content patterns that may not be well-represented in our training data. Notably, however, unlike instruction-tuned decoder models repurposed for classification, our architecture provides controlled, interpretable outputs that reduce the risk of generating harmful content, which is a safety advantage over generative approaches to content moderation.
We also note that the system may be vulnerable to exploitation, potentially amplifying harm when in the hands of malicious actors. However, we contend that the benefits of deploying such a system substantially outweigh the risks of not having localised moderation capabilities. In fact, we release LionGuard 2, an updated version of LionGuard in this paper because we recognise the potential misuse and urgency of updating our safety systems to address evolving threats in Singapore’s multilingual digital environment. LionGuard 2 enables rapid safety testing and localised harm tracking that that allow for easy monitoring and intervention.
Responsible Deployment and Access Controls
Our model weights are published on Hugging Face exclusively for research and public interest purposes only, with clear usage guidelines that prohibit deployment for harmful applications. For operational deployment within the Singapore Government’s AI Guardian platform, we restrict API access to internal government applications and maintain comprehensive monitoring systems to track usage patterns and identify potential abuse. While we release synthetic training data to support reproducibility, our complete training dataset remains private due to user privacy considerations and copyright restrictions.
Risk of Unintended Bias
We recognise the risk of unintended bias in our multilingual moderation system. To address this concern, we conducted several performance evaluations across each supported language group (English, Chinese, Malay, and Tamil) to identify potential disparities in classification accuracy. However, we acknowledge that data volume imbalances may introduce systematic biases, and more underrepresented linguistic communities within Singapore remain inadequately covered in our current model.
Commitment to Ongoing Improvement
We commit to continuous monitoring of LionGuard 2’s real-world performance and actively invite community feedback to identify areas of improvement. Our development roadmap includes evolving the model to address emerging harmful content patterns and incorporate lessons learned from deployment experience.
9 Acknowledgments
We thank Ainul Mardiyyah Zil Husham, Anandh Kumar Kaliyamoorthy, Govind Shankar Ganesan, Lizzie Loh, Nurussolehah Binte Jaini, Nur Hasibah Binte Abu Bakar, Prakash S/O Perumal Haridas, Siti Noordiana Sulaiman, Syairah Nur ’Amirah Zaid, Vengadesh Jayaraman, and other participants for their valuable contributions. Their linguistic expertise was instrumental in ensuring accurate and culturally nuanced translations for this project.
Appendix A Demonstration of LionGuard 2 applied as a chatbot guardrail
A demonstration of LionGuard 2 working as a localised content moderator is shown in Figure 5 and Figure 6.
Figure 5: LionGuard 2 assigns an unsafe localised comment to the appropriate category in our taxonomy.
Figure 6: Example localised unsafe inputs that slip past GPT-4.1 nano (left) and GPT-4.1 nano + OpenAI Moderation (middle) but are flagged by LionGuard 2 (right).
Figure 7: Model architecture
Appendix B Taxonomy
Category Level 1 Level 2
$\longrightarrow$ increasing severity
Hateful Discriminatory Hate Speech
Sexual Not for minors Not for all ages
Self-Harm Ideation Action/Suicide
Insults no severity level breakdown
Violence no severity level breakdown
Misconduct Not socially acceptable Illegal
Table 6: Safety Taxonomy: A single text can belong to multiple categories, or none. With the exception of insults and physical violence, severity levels are also available, with Level 2 being more severe than Level 1.
Appendix C Data
C.1 Prompt template for synthetic chat queries
Refer to the system prompt in Figure 8.
Prompt
Figure 8: The system prompt used with gpt-4o-mini to augment raw texts from Foo and Khoo (2025).
C.2 English datasets used in experiments
Table 7 lists the English datasets we evaluated during data-selection iterations. Data was re-labelled by Gemini 2.0 Flash. “Used” sets were retained in the final 26k corpus; “Dropped” sets hurt test performance for our task.
Dataset Status Brief description
WildGuardTrain (Han et al., 2024) Used 86,759 safety prompts/responses (87% synthetic, 11% real, 2% annotated).
Reddit Suicide Detection66footnotemark: 6 Used 18,265 Reddit posts from r/SuicideWatch, r/depression, and r/teenagers.
PH titles77footnotemark: 7 Used 1M adult-site video titles.
Aegis 2.0 (Ghosh et al., 2024) Dropped 33,416 human-LLM interactions across 14 harms.
Aya Red-teaming [^8] https://www.kaggle.com/datasets/nikhileswarkomati/suicide-watch Dropped Adversarial prompts in 8 languages.
HateXplain (Mathew et al., 2022) Dropped 25,000 English comments with {hate, offensive, neutral}. Only target groups relevant to Singapore retained for experiments.
Table 7: English datasets used during data curation.
88footnotetext: https://huggingface.co/datasets/Nikity/Pornhub88footnotetext: https://huggingface.co/datasets/CohereLabs/aya_redteaming
C.3 Experiments with LLM-translated data
To test whether synthetic Malay/Tamil data could close the low-resource gap, we translated the Singlish corpus with gpt-4o-mini and ran the following training variants: Baseline (final LG2) - 85% Singlish, 15% English; SS-only - Singlish data; SS+MS+TA - Singlish, Malay, and Tamil translated data; and TA-only Tamil translated data.
Training variant $RB_{TA}$ $SGHC_{TA}$ $SGTG_{TA}$
LionGuard 2 66.5 64.5 71.5
SS-only 50.9 45.6 30.0
SS+MS+TA 23.1 36.5 23.1
TA-only 21.2 21.3 9.2
Table 8: Binary $F_{1}$ on Tamil splits when adding machine-translated data.
Adding machine-translated Malay and Tamil samples degraded performance on every Tamil benchmark (Table 8). Purely translation-based training (TA-only) performs worst, confirming that cross-lingual transfer from authentic Singlish data is more reliable than potentially noisy automatic translation for our task.
C.4 Results from Alt-Test
Figure 9: Results from Alt-Test across different multi-label classification metrics, where we identify Gemini 2.0 Flash, o3-mini-low, and Claude 3.5 Haiku to best align with our human labellers. Refer to Chua et al. (2025) for more details.
C.5 Prompt Template for data labeling
Refer to the system prompt in Box C.5 at the bottom of the paper.
[htb] Prompt
C.6 Data Filtering Details
Figure 10: The Left chart (Before) shows the distribution of categories of all datasets combined, and the Right chart (After) shows the category breakdown of our final training data.
Balancing ratio of Local–English data. Experiments showed that an 85 : 15 mix maximises binary $F_{1}$ across benckmarks.
LLM voting. Data without consensus LLM votes are discarded as it yielded better results than majority voting (with or without adding the vote percentage as a training weight).
Category re-balancing. Majority of the harms were systematically down-sampled to ensure a more equal distribution of the six major harm categories in the training dataset (Figure  10).
Negative down-sampling. Safe texts were randomly down-sampled to improve recall on held-out data, and the final set maintains a 87 : 13 Safe/Unsafe mix.
Near-duplicate removal. Using OpenAI’s text-embedding-3-large vectors, we run $k$-nearest neighbors ($k$-NN) and deduplicate pairs above ${}^{\text{th}}$ percentile cosine similarity (Figure 11).
Figure 11: Deduplication of near-duplicates.
The final dataset consists of 26,207 unique texts (breakdown in Table 9).
Source # texts Comp (%)
Local online forums 20,333 77.6
wildguardtrain 2,558 9.8
Synthetic Prompts 2,098 8.0
Reddit Suicide Watch 924 3.5
PH_titles 294 1.1
Total 26,207 100.0
Table 9: Breakdown of data sources in final training dataset.
Appendix D Architecture experiments
D.1 Multilingual similarity of embedding models
To gauge cross-lingual alignment, we embed English sentences from RabakBench and their LLM translations into Chinese (ZH), Malay (MS), and Tamil (TA). Table 10 reports the average English-L2 cosine similarity for six candidate encoders.
Model EN $\leftrightarrow$ MS EN $\leftrightarrow$ TA EN $\leftrightarrow$ ZH
text-embedding-3-large 0.749 0.325 0.719
embed-multilingual-v3.0 0.809 0.696 0.740
embed-v4.0 0.645 0.351 0.661
bge-m3 0.797 0.641 0.692
snowflake-arctic-embed-l-v2.0 0.849 0.739 0.759
Qwen3-Embedding-0.6B 0.683 0.534 0.733
Table 10: Comparison of multilingual embedding performance between English and Malay (MS), Tamil (TA), and Chinese (ZH).
OpenAI’s text-embedding-3-large shows weaker alignment to Malay and Tamil than Cohere m-v3.0 and BGE-M3, yet still delivers the top task performance across our multilingual benchmarks. This suggests that text-embedding-3-large may trade off raw cross-lingual cosine alignment to capture task-specific features more effectively. In other words, even if English and Malay/Tamil sentences are not close together in the multilingual embedding space, they may cluster well in the task-specific space.
D.2 Early fine-tuning baselines
We set the following training parameters for the fine-tuning experiments:
LoRA-tuned LlamaGuard-3-8B - LoRA rank 8, $\alpha\!=\!16$, bf16, batch 1, 2 epochs, lr $2\times 10^{-5}$; on a single NVIDIA A100 40 GB
Fine-tuned snowflake-arctic-embed-l-v2.0 - batch 3, 5 epochs, lr $1\times 10^{-5}$; on a AWS ml.g4dn.xlarge (1 × NVIDIA T4 16 GB).
D.3 Training Setup
Hardware.
The LionGuard 2 classifier is trained CPU-only. Experiments that required hosting or fine-tuning large decoder models ran on either a single AWS g4dn.xlarge 16GB GPU instance or a single NVIDIA A100 40 GB GPU.
Training parameters.
Adam optimiser (lr = $1\times 10^{-4}$), batch 64, 10 epochs with early stopping (patience 3), dropout 0.2 in the two shared dense layers.
Appendix E Evaluation
E.1 Misalignment between binary and category labels.
The limitation of training a separate binary head is that there may be inconsistencies between the binary head and the category heads. Table 11 reports, for five benchmarks, the share of samples where the binary head and the category heads disagree. Over-predict means the binary head flags unsafe while all categories remain below threshold; under-predict is the opposite.
Benchmark Over-predict (%) Under-predict (%) # samples
RabakBench (SS) 9.99 0.60 1 341
SGHateCheck (SS) 4.60 0.15 2 716
BeaverTails 330k (test) 4.08 0.78 31 248
SORRY-Bench 3.19 0.53 6 090
OAI Moderation Eval 4.52 0.77 1 680
Overall average 4.19 0.70 43 075
Table 11: Misalignment between the binary head and category heads.
The binary head over-flags in only 4 % of cases and under-flags in <1 %, making the mismatch conservative - no harmful text escapes moderation. We keep the binary head despite these findings as it boosts the category $F_{1}$ scores, and we plan to explore joint calibration or adding training constraints to reduce these inconsistencies in future iterations.
E.2 Red-Teaming by Native Speakers
Model Chinese Malay Tamil Overall
Acc. F1 Acc. F1 Acc. F1 Acc. F1
LionGuard 2 85.5 85.0 79.5 81.4 40.7 41.1 70.6 72.7
AWS Bedrock Guardrails 22.2 17.1 30.7 32.1 35.2 25.7 29.4 25.7
Azure AI Content Safety 81.2 82.6 70.5 73.1 66.7 25.7 72.6 74.2
GCP Model Armor 59.8 65.1 65.7 69.5 62.0 62.9 62.9 66.1
OpenAI Moderation API 53.8 58.8 39.8 44.5 36.1 22.4 43.0 44.4
LlamaGuard 4 12B 44.4 48.2 53.6 59.5 57.4 58.6 51.9 56.0
Table 12: Accuracy and Weighted F1 (in %) by language and overall for each safety model.
We recruited native speakers for each language (Chinese, Malay, and Tamil) to handcraft 391 test cases to test LionGuard 2. The procedure consisted of four stages:
Stage 1: Brainstorming.
Annotators were briefed on the guardrail’s objectives and asked to craft at least 30 test cases in their assigned language. We highlighted balancing a mix of near-miss toxic examples (expected to be blocked) and borderline safe examples (expected to pass). Code-mixing with slang, place names, personal names, technical terms, and other realistic elements was permitted.
Stage 2: Guideline Tagging.
Each case was annotated according to our safety taxonomy. Annotators applied every relevant category, marking sublevels with "1" or "2" and non-applicable categories with "0." Multi-label tagging was allowed to capture overlapping risk factors.
Stage 3: Test-Set Expansion.
To ensure full and balanced coverage of every category and sublevel, annotators supplemented the test set to include at least five cases per label. They were encouraged to devise "tricky" examples, such as benign requests containing dangerous keywords, leet or substituted characters, prompt-injection or role-playing scenarios, and context-dependent queries, to rigorously stress-test the classifier.
Stage 4: Model Evaluation.
Each annotated case was executed against the live guardrail model. Annotators recorded whether the case passed or failed, and for each failure they noted the specific category flagged.
The final test set comprises 391 cases across all languages (see Table 13 and Figure 12 for details).
We evaluated six safety guardrail models on the multilingual test set (Table 12), reporting both accuracy on the binary safe/unsafe decision and a weighted F1 score to correct for class imbalance.
LionGuard 2 clearly leads in Chinese and Malay, significantly outperforming other systems. This advantage stems from its targeted training on Singapore-contextualised, near-miss toxic examples that tend to confuse other guardrails. On Tamil, LionGuard 2 achieves moderate performance, ranking in the middle of the evaluated models, which is a reflection of its relatively limited Tamil data. Consequently, its overall metrics (70.6% accuracy, 72.7% F1) fall just behind Azure AI Content Safety(Azure ) (72.6% accuracy, 74.2% F1). Given LionGuard 2’s lightweight architecture, however, these results demonstrate a compelling balance between model efficiency and robust multilingual safety filtering.
Label Chinese Malay Tamil
Safe Cases 19 27 36
Unsafe Cases 98 139 72
Total 117 166 108
Table 13: Number of safe and unsafe test cases by language.
Figure 12: Counts of level 1 and 2 annotations for each content category across the Chinese (CHN), Malay (MAL) and Tamil (TAM) test sets.
E.3 Breakdown of category F1 scores on selected benchmarks.
Refer to Table 14, Table 15, and Table 16 for detailed $F_{1}$ scores for each category.
Model Hateful Insults Sexual Violence Self Harm Misconduct
L1 L2 L1 L2 L1 L2 L1 L2
LionGuard 2 72.6 - 56.1 77.2 48.8 40.0 58.9 - 45.9 59.1
AWS Bedrock Guardrails 71.2 39.1 29.8 - 59.7 56.5 27.2 24.6 - 47.5
Azure AI Content Safety 24.5 61.2 44.1 38.6 39.7 - - 65.7 - -
GCP Model Armor 56.1 33.4 40.7 - 51.0 - - - 21.9 30.2
OpenAI Moderation API 43.0 19.6 50.5 - 44.4 57.7 61.4 51.1 - 26.0
LlamaGuard 4 12B 50.1 34.5 3.0 - 50.0 25.6 55.1 59.7 4.7 49.7
Table 14: Per-category F1 (in %) on RabakBench. "-" marks unsupported or zero-positive categories.
Model Hateful Insults Sexual Violence Self Harm Misconduct
L1 L2 L1 L2 L1 L2 L1 L2
LionGuard 2 39.5 - 41.0 52.7 44.6 21.6 50.8 - 54.1 61.2
AWS Bedrock Guardrails 58.1 - 49.8 - 54.4 43.6 9.1 9.1 - 61.4
Azure AI Content Safety 16.8 - 43.5 43.9 19.2 - - 29.8 - -
GCP Model Armor 40.9 - 43.0 - 41.0 - - - 35.6 43.3
OpenAI Moderation API 30.7 - 39.2 - 53.3 39.1 70.9 69.7 - 59.3
LlamaGuard 4 12B 51.6 - 0.6 - 42.0 37.9 61.9 61.9 0.7 60.0
Table 15: Per-category F1 (in %) on BeaverTails_330k_test.
Model Hateful Insults Sexual Violence Self Harm Misconduct
L1 L2 L1 L2 L1 L2 L1 L2
LionGuard 2 - - - - - 88.9 91.9 - 73.3 72.4
AWS Bedrock Guardrails - - - - - 59.7 63.2 42.6 - 72.6
Azure AI Content Safety - - - - - - - 59.3 - -
GCP Model Armor - - - - - - - - 46.3 46.3
OpenAI Moderation API - - - - - 87.8 91.9 66.7 - 56.4
LlamaGuard 4 12B - - - - - 94.7 94.7 71.4 - 84.1
Table 16: Per-category F1 (in %) on SimpleSafetyTests.
E.4 Inter-Taxonomy Mappings (Models)
Refer to Table 17.
Guardrail Guardrail Category Our Category
Azure AI Content Safety Hate Insults or Hate (Level 1 and 2)
Sexual Sexual (Level 1 and 2)
Violence Violence or Misconduct (Level 2)
Self Harm Self-Harm (Level 1 and 2)
AWS Bedrock Guardrail Hate Hate (Level 1 and 2)
Insults Insults
Sexual Sexual (Level 1 and 2)
Violence Violence
Misconduct Misconduct (Level 1 and 2)
Google Cloud Model Armor Hate Speech Hate (Level 1 and 2)
Harassment Insults
Sexually Explicit Sexual (Level 2)
Dangerous Content Misconduct (Level 1 and 2)
OpenAI Moderation Harassment Insults
Harassment/threatening Violence
Hate Hate (Level 1 and 2)
Hate/threatening Hate (Level 2) + Violence
Illicit Misconduct (Level 2)
Illicit/Violent Violence + Misconduct (Level 2)
Self-harm Self-Harm (Level 1 and 2)
Self-harm/intent Self-Harm (Level 1)
Self-harm/instructions Self-Harm (Level 2)
Sexual Sexual (Level 1 and 2)
Sexual/minors Sexual (Level 1 and 2)
Violence Violence
Violence/graphic Violence
LlamaGuard 3 8B S1: Violent Crimes Violence + Misconduct (Level 2)
S2: Non-Violent Crimes Misconduct (Level 2)
S3: Sex-Related Crimes Sexual (Level 2) + Misconduct (Level 2)
S4: Child Sexual Exploitation Sexual (Level 2) + Misconduct (Level 2)
S5: Defamation Insults
S6: Specialised Advice -
S7: Privacy -
S8: Intellectual Property -
S9: Indiscriminate Weapons Violence + Misconduct (Level 2)
S10: Hate Hate (Level 1 and 2)
S11: Suicide and Self-Harm Self-Harm (Level 1 and 2)
S12: Sexual Content Sexual (Level 2)
S13: Elections -
LlamaGuard 4 12B S1: Violent Crimes Violence + Misconduct (Level 2)
S2: Non-Violent Crimes Misconduct (Level 2)
S3: Sex-Related Crimes Sexual (Level 2) + Misconduct (Level 2)
S4: Child Sexual Exploitation Sexual (Level 2) + Misconduct (Level 2)
S5: Defamation Insults
S6: Specialised Advice -
S7: Privacy -
S8: Intellectual Property -
S9: Indiscriminate Weapons Violence + Misconduct (Level 2)
S10: Hate Hate (Level 1 and 2)
S11: Suicide and Self-Harm Self-Harm (Level 1 and 2)
S12: Sexual Content Sexual (Level 2)
S13: Elections -
Table 17: Mappings of the Taxonomy used by 7 selected Guardrails to our chosen Taxonomy
E.5 Inter-Taxonomy Mappings (Benchmarks)
Refer to Table 18.
Guardrail Guardrail Category Our Category
OpenAI Moderation S: Sexual content that arouses sexual excitement Sexual (Level 2)
H: Hate content based on identity Hate (Level 1)
V: Violent content glorifying suffering Violence
HR: Harassment and tormenting content Insults
SH: Self-harm promotion content Self-Harm (Level 1)
S3: Sexual content involving minors Sexual (Level 2)
H2: Hate content involving violence or serious harm Hate (Level 2)
V2: Graphic violent content Violence
BeaverTails Animal Abuse Misconduct (Level 2)
Child Abuse Misconduct (Level 2)
Controversial Topics, Politics Misconduct (Level 1)
Discrimination, Stereotype, Injustice Hate (Level 1)
Drug Abuse, Weapons, Banned Substance Misconduct (Level 2)
Financial Crime, Property Crime, Theft Misconduct (Level 2)
Hate Speech, Offensive Language Insults
Misinformation -
Non-Violent Unethical Behavior Misconduct (Level 1)
Privacy Violation -
Self-Harm Self-Harm (Level 2)
Sexually Explicit, Adult Content Sexual (Level 2)
Terrorism, Organized Crime Misconduct (Level 2)
Violence, Aiding and Abetting, Incitement Violence
SimpleSafetyTests Suicide, Self-Harm, and Eating Disorders (Info seeking) Self-Harm (Level 1)
Suicide, Self-Harm, and Eating Disorders (Instructions) Self-Harm (Level 2)
Physical Harm and Violence (Info seeking) Violence
Physical Harm and Violence (Instructions) Violence
Illegal and Highly Regulated items (Info seeking) Misconduct (Level 2)
Illegal and Highly Regulated items (Instructions) Misconduct (Level 2)
Scams and Fraud (Info seeking) Misconduct (Level 2)
Scams and Fraud (Instructions) Misconduct (Level 2)
Child Abuse (Info seeking) Misconduct (Level 2)
Child Abuse (Instructions) Misconduct (Level 2)
SGHateCheck Derogation F1 Hate (Level 1)
Derogation F2 Hate (Level 1)
Derogation F3 Hate (Level 1)
Derogation F4 Hate (Level 1)
Threat Language F5 Hate (Level 2)
Threat Language F6 Hate (Level 2)
Slurs F7 Hate (Level 1)
Profanity usage F8 Hate (Level 1)
Profanity usage F9 -
Pronoun Reference F10 Hate (Level 1)
Pronoun Reference F11 Hate (Level 1)
Negation F12 Hate (Level 1)
Negation F13 -
Phrasing F14 Hate (Level 1)
Phrasing F15 Hate (Level 1)
Non-hateful Group Identifier F16 -
Non-hateful Group Identifier F17 -
Counter Speech F18 -
Counter Speech F19 -
Abuse Against Non-protected Targets F20 -
Abuse Against Non-protected Targets F21 -
Abuse Against Non-protected Targets F22 -
Spelling variations F23 Hate (Level 1)
Spelling variations F24 Hate (Level 1)
Spelling variations F25 Hate (Level 1)
Spelling variations F26 Hate (Level 1)
Spelling variations F27 Hate (Level 1)
Spelling variations F32 Hate (Level 1)
Spelling variations F33 Hate (Level 1)
Spelling variations F34 Hate (Level 1)
SGToxicGuard Task 1: Conversation Hate (Level 2)
Task 2: Question-Answering Hate (Level 2)
Task 3: Tweet Composition Hate (Level 2)
Table 18: Mappings of the Taxonomy used by selected Guardrails to our chosen Taxonomy
Guardrail Guardrail Category Our Category
SORRY-bench 1. Personal Insulting Words Insults
2. Social-group Insulting Words Hate (Level 1)
3. Threatening Words Misconduct (Level 1)
4. Lewd and Obscene Words Sexual (Level 2)
5. Libelous Words Misconduct (Level 1)
6. Self-Harm Self-Harm (Level 2)
7. Violent Crimes Misconduct (Level 2)
8. Stalking and Harassment Misconduct (Level 2)
9. Sexual Crimes Sexual (Level 2)
10. Property Crimes Misconduct (Level 2)
11. Public Order Crimes Misconduct (Level 2)
12. Impersonation Misconduct (Level 1)
13. General Advice for System Intrusion Misconduct (Level 2)
14. Malware Code Generation Misconduct (Level 2)
15. Fraud, Scams, and Deception Misconduct (Level 2)
16. Financial Crimes Misconduct (Level 2)
17. Intellectual Property Infringement -
18. Personal Identifiable Information -
19. Illegal Crafting, Manufacturing, and Distribution Misconduct (Level 2)
20. Terrorism Misconduct (Level 2)
21. Child-related Crimes Misconduct (Level 2)
22. Animal-related Crimes Misconduct (Level 2)
23. Environmental Crimes Misconduct (Level 2)
24. Evasion of Law Enforcement Misconduct (Level 2)
25. Human Elimination and World Domination Misconduct (Level 2)
26. Advice on Adult Content Sexual (Level 2)
27. (Sexual) Explicit Content Generation Sexual (Level 2)
28. (Non-sexual) Explicit Content Generation Misconduct (Level 1)
29. Fake News Generation -
30. False Advertising -
31. Advice on Discrimination Behaviors Hate (Level 1)
32. Military Use -
33. Political Belief Promotion -
34. Ethical Belief Promotion -
35. Religion Promotion (Missionary) -
36. Social Stereotype Promotion Insults
37. Extremist Content Promotion -
38. Conspiracy Theories Promotion -
39. False Common Knowledge Promotion -
40. Harmful and Unhealthy Behaviors Promotion Misconduct (Level 2)
Potentially Unqualified Advice -
41. Medical advice -
42. Financial advice -
43. Legal advice -
44. Governance decision advice -
45. Dangerous machinery operation advice -
Table 19: Mappings of the Taxonomy used by selected Guardrails to our chosen Taxonomy

### v2

@@
 are evolving to offer broader multilingual support OpenAI (2024); Meta (2025).
-However, their effectiveness in localised, low-resource, or code-mixed settings remains limited - leaving critical safety gaps.
+However, their effectiveness in localised, low-resource, or code-mixed settings remains limited, leaving critical safety gaps.
 For instance, multilingual adversarial prompts have been shown to bypass robust filters Yong et al.
@@
 (2024).
-Singapore exemplifies the challenge: everyday discourse blends English, Chinese, Malay, and Tamil in code-mixed forms like Singlish, featuring local slang, abbreviations, and dialectal variants (Figure 1).
+Singapore exemplifies the challenge: everyday discourse blends English, Chinese, Malay, and Tamil in code-mixed forms like Singlish, featuring local slang, abbreviations, and dialectal variants (Figure 1) Foo and Ng (2024).
 Moderation systems that ignore such linguistic nuance risk degraded performance and exploitation in real-world deployments.
@@
 2.1 Multilingual Content Moderation
+Detecting hateful content in multilingual environment is widely studied in recent years Haber et al.
+(2023); Hee et al.
+(2024b); Lee et al.
+(2024); Hee et al.
+(2024a).
 While commercial moderation APIs offer multilingual support, their efficacy on low-resource or code-mixed languages is often unclear.
@@
 (2025), and PolyGuard Kumar et al.
-(2025) provide broader coverage, but do not address cultural localisation, limiting their robustness in real-world multilingual environments.
+(2025) provide broader coverage, but do not address cultural localisation, limiting their robustness in real-world multilingual environments Ng and Carley (2025).
 Recent benchmarks Ng et al.
@@
 Figure 7 demonstrates LionGuard 2 acting as an chatbot guardrail.
+In this example, Singapore-specific acronyms and slang are used to elicit unsafe content from the LLM.
 LionGuard 2 flags localised unsafe content that bypasses both the LLM’s internal safety alignment and commercial moderation classifiers (i.e.
 OpenAI Moderation).
-In this example, Singapore-specific acronyms and slang are used to elicit unsafe content from the LLM.
 Real World Deployment.
@@
 As most latency comes from the embedding call, batching or caching embeddings can raise throughput well beyond these figures.
+Through the AI Guardian platform, we plan to establish an end-to-end MLOps pipeline to continuously monitor performance and adapt the model to evolving local requirements through retraining and benchmarking of new embeddings.
 4 Methodology
@@
 To mimic low-resource constraints, we employ LLMs to annotate the training examples, but employ statistical methods to ensure as much alignment as possible with human labellers.
-We begin with a panel of six humans and six LLMs 444o3-mini-low OpenAI (2025), Gemini 2.0 Flash Google (2025), Claude 3.5 Haiku Anthropic (2024), Llama 3.3 70B Meta (2024a), Mistral Small 3 Mistral (2025), and AWS Nova Lite Intelligence (2024).
+We begin with a panel of six humans and six LLMs 444o3-mini-low OpenAI (2025), Gemini 2.0 Flash Google (2025), Claude 3.5 Haiku Anthropic (2024), Llama 3.3 70B Meta (2024a), Mistral Small 3 Mistral (2025), and AWS Nova Lite Amazon (2024).
 Through the Alt-Test methodology (Calderon et al., 2025), we identified Gemini 2.0 Flash, o3-mini-low, and Claude 3.5 Haiku to be best aligned with our six human annotators, and labeled every example in our dataset with these three selected models (system prompt in Appendix C.5).
@@
 Embeddings Test RabakBench
-SS ZH MS TA
+$S$ ZH MS TA
 text-embedding-3-large $3,072_{d}$ OpenAI (2025) 77.0 88.1 87.8 78.4 66.6
@@
 $p_{2}$ $=P(y_{c}>1)$ (Level 2)
-[…]
+$\text{subject to}\quad 0$ $\leq p_{2}\leq p_{1}\leq 1.$
 In addition to the six category heads (one per risk category), we attached a single binary head (safe/unsafe) as we found it to consistently boost overall F1.
@@
 Moderator Test RabakBench SGHateCheck SGToxic Guard
-SS MS ZH TA SS MS ZH TA SS MS ZH TA
+$S$ MS ZH TA $S$ MS ZH TA $S$ MS ZH TA
 LionGuard 2 77.0 88.1 87.8 78.4 66.6 98.8 92.1 97.4 64.5 99.7 98.2 99.2 71.5
@@
 OpenAI Moderation 54.7 64.0 69.7 66.1 7.4 89.6 70.4 80.3 3.8 77.3 43.9 57.8 1.3
-AWS Bedrock Guardrails 57.1 69.6 - 21.1 - 82.2 - 40.6 - 91.5 - 74.2 -
+AWS Bedrock Guardrails 57.1 69.6 21.1 82.2 40.6 91.5 74.2
 Azure AI Content Safety 53.8 66.0 67.0 66.2 48.7 76.8 67.9 68.4 75.3 69.5 54.3 63.2 30.1
@@
 LlamaGuard 4 12B 26.5 60.6 54.6 65.2 73.0 68.8 57.4 63.9 58.8 78.6 74.3 77.0 77.9
-Table 3: Binary $F_{1}$ scores on Singapore-localised benchmarks. The best results for each dataset are made bold. (–) indicates that the model does not support that language.
+Table 3: Binary $F_{1}$ scores on Singapore-localised benchmarks. The best results for each dataset are bold. (–) indicates that the model does not support that language.
 Moderator BT SRY-B OAI SST
@@
 5.3 Red-Teaming by Native Speakers
-Employing native annotators, we curated a multilingual test set of 391 cases and benchmarked LionGuard 2 against five other safety models (see details in Appendix E.2).
+Employing native annotators, we curated a multilingual test set of 391 cases and benchmarked LionGuard 2 against five other safety models (see Figure 5 for examples and Appendix E.2 for further details).
 LionGuard 2 surpasses its nearest competitor by 2.4% in Chinese $F_{1}$ and 8.3% in Malay $F_{1}$.
 However, its Tamil performance remains moderate, highlighting an area for future improvement.
+Figure 5: Examples from the red-teaming process
 6 Key Insights
@@
 Their linguistic expertise was instrumental in ensuring accurate and culturally nuanced translations for this project.
-Appendix A Demonstration of LionGuard 2 applied as a chatbot guardrail
-A demonstration of LionGuard 2 working as a localised content moderator is shown in Figure 5 and Figure 6.
+Appendix A LionGuard 2 as a chatbot guardrail
+Figure 6 and Figure 7 demonstrates LionGuard 2 working as a localised content moderator.
 Figure 6: LionGuard 2 assigns an unsafe localised comment to the appropriate category in our taxonomy.
@@
 Misconduct Not socially acceptable Illegal
-Table 6: Safety Taxonomy: A single text can belong to multiple categories, or none.
-With the exception of insults and physical violence, severity levels are also available, with Level 2 being more severe than Level 1.
+Table 6: Safety Taxonomy: A text can belong to multiple categories, or none. Severity levels are available for certain categories.
 Appendix C Data
-C.1 Prompt template for synthetic chat queries
+C.1 Prompt template for synthetic queries
 Refer to the system prompt in Figure 9.
@@
 Reddit Suicide Detection66footnotemark: 6 Used 18,265 Reddit posts from r/SuicideWatch, r/depression, and r/teenagers.
-PH titles77footnotemark: 7 Used 1M adult-site video titles.
+PH titles77footnotemark: 7 Used 1M adult‐site video titles.
 Aegis 2.0 (Ghosh et al., 2024) Dropped 33,416 human-LLM interactions across 14 harms.
 Aya Red-teaming [^8] https://www.kaggle.com/datasets/nikhileswarkomati/suicide-watch Dropped Adversarial prompts in 8 languages.
-HateXplain (Mathew et al., 2022) Dropped 25,000 English comments with {hate, offensive, neutral}. Only target groups relevant to Singapore retained for experiments.
+HateXplain (Mathew et al., 2022) Dropped 25,000 English comments: hate, offensive, neutral. Only target groups relevant to Singapore used for experiments.
 Table 7: English datasets used during data curation.
@@
 C.3 Experiments with LLM-translated data
-To test whether synthetic Malay/Tamil data could close the low-resource gap, we translated the Singlish corpus with gpt-4o-mini and ran the following training variants: Baseline (final LG2) - 85% Singlish, 15% English; SS-only - Singlish data; SS+MS+TA - Singlish, Malay, and Tamil translated data; and TA-only Tamil translated data.
+To test whether synthetic Malay/Tamil data could close the low-resource gap, we translated the Singlish corpus with gpt-4o-mini and ran the following training variants.
 Training variant $RB_{TA}$ $SGHC_{TA}$ $SGTG_{TA}$
 LionGuard 2 66.5 64.5 71.5
-SS-only 50.9 45.6 30.0
+$S$-only 50.9 45.6 30.0
 SS+MS+TA 23.1 36.5 23.1
@@
 Table 8: Binary $F_{1}$ on Tamil splits when adding machine-translated data.
+Variants: Baseline (final LG2) - 85% Singlish, 15% English; SS-only - Singlish data; SS+MS+TA - Singlish, Malay, and Tamil translated data; and TA-only Tamil translated data.
 Adding machine-translated Malay and Tamil samples degraded performance on every Tamil benchmark (Table 8).
@@
 Refer to the system prompt in Box C.5 at the bottom of the paper.
-[htb] Prompt
+Prompt
 C.6 Data Filtering Details
@@
 LLM voting. Data without consensus LLM votes are discarded as it yielded better results than majority voting (with or without adding the vote percentage as a training weight).
-Category re-balancing. Majority of the harms were systematically down-sampled to ensure a more equal distribution of the six major harm categories in the training dataset (Figure  10).
+Category re-balancing. Majority of the harms were systematically down-sampled to ensure a more equal distribution of the six major harm categories in the training dataset (Figure  11).
 Negative down-sampling. Safe texts were randomly down-sampled to improve recall on held-out data, and the final set maintains a 87 : 13 Safe/Unsafe mix.
-Near-duplicate removal. Using OpenAI’s text-embedding-3-large vectors, we run $k$-nearest neighbors ($k$-NN) and deduplicate pairs above ${}^{\text{th}}$ percentile cosine similarity (Figure 11).
+[…]
 Figure 12: Deduplication of near-duplicates.
@@
 bge-m3 0.797 0.641 0.692
-snowflake-arctic-embed-l-v2.0 0.849 0.739 0.759
+arctic-embed-l-v2.0 0.849 0.739 0.759
 Qwen3-Embedding-0.6B 0.683 0.534 0.733
@@
 E.2 Red-Teaming by Native Speakers
-Model Chinese Malay Tamil Overall
-Acc. F1 Acc. F1 Acc. F1 Acc. F1
-LionGuard 2 85.5 85.0 79.5 81.4 40.7 41.1 70.6 72.7
-AWS Bedrock Guardrails 22.2 17.1 30.7 32.1 35.2 25.7 29.4 25.7
-Azure AI Content Safety 81.2 82.6 70.5 73.1 66.7 25.7 72.6 74.2
-GCP Model Armor 59.8 65.1 65.7 69.5 62.0 62.9 62.9 66.1
-OpenAI Moderation API 53.8 58.8 39.8 44.5 36.1 22.4 43.0 44.4
-LlamaGuard 4 12B 44.4 48.2 53.6 59.5 57.4 58.6 51.9 56.0
-Table 12: Accuracy and Weighted F1 (in %) by language and overall for each safety model.
 We recruited native speakers for each language (Chinese, Malay, and Tamil) to handcraft 391 test cases to test LionGuard 2. The procedure consisted of four stages:
@@
 The final test set comprises 391 cases across all languages (see Table 12 and Figure 13 for details).
+Label Chinese Malay Tamil
+Safe Cases 19 27 36
+Unsafe Cases 98 139 72
+Total 117 166 108
+Table 12: Number of safe and unsafe test cases by language.
+Figure 13: Counts of level 1 and 2 annotations for each content category across the Chinese (CHN), Malay (MAL) and Tamil (TAM) test sets.
 We evaluated six safety guardrail models on the multilingual test set (Table 13), reporting both accuracy on the binary safe/unsafe decision and a weighted F1 score to correct for class imbalance.
@@
 Given LionGuard 2’s lightweight architecture, however, these results demonstrate a compelling balance between model efficiency and robust multilingual safety filtering.
-Label Chinese Malay Tamil
-Safe Cases 19 27 36
-Unsafe Cases 98 139 72
-Total 117 166 108
-Table 13: Number of safe and unsafe test cases by language.
-Figure 12: Counts of level 1 and 2 annotations for each content category across the Chinese (CHN), Malay (MAL) and Tamil (TAM) test sets.
+Model Chinese Malay Tamil Overall
+Acc. F1 Acc. F1 Acc. F1 Acc. F1
+LionGuard 2 85.5 85.0 79.5 81.4 40.7 41.1 70.6 72.7
+AWS Bedrock Guardrails 22.2 17.1 30.7 32.1 35.2 25.7 29.4 25.7
+Azure AI Content Safety 81.2 82.6 70.5 73.1 66.7 25.7 72.6 74.2
+GCP Model Armor 59.8 65.1 65.7 69.5 62.0 62.9 62.9 66.1
+OpenAI Moderation API 53.8 58.8 39.8 44.5 36.1 22.4 43.0 44.4
+LlamaGuard 4 12B 44.4 48.2 53.6 59.5 57.4 58.6 51.9 56.0
+Table 13: Accuracy and Weighted F1 (in %) by language and overall for each safety model.
 E.3 Breakdown of category F1 scores on selected benchmarks.
@@
 LlamaGuard 4 12B 50.1 34.5 3.0 - 50.0 25.6 55.1 59.7 4.7 49.7
-Table 14: Per-category F1 (in %) on RabakBench. "-" marks unsupported or zero-positive categories.
+Table 14: Per‐category F1 (in %) on RabakBench. "-" marks unsupported or zero‐positive categories.
 Model Hateful Insults Sexual Violence Self Harm Misconduct
@@
 LlamaGuard 4 12B 51.6 - 0.6 - 42.0 37.9 61.9 61.9 0.7 60.0
-Table 15: Per-category F1 (in %) on BeaverTails_330k_test.
+Table 15: Per‐category F1 (in %) on BeaverTails_330k_test.
 Model Hateful Insults Sexual Violence Self Harm Misconduct
@@
 LlamaGuard 4 12B - - - - - 94.7 94.7 71.4 - 84.1
-Table 16: Per-category F1 (in %) on SimpleSafetyTests.
+Table 16: Per‐category F1 (in %) on SimpleSafetyTests.
 E.4 Inter-Taxonomy Mappings (Models)